OpenStudy (anonymous):

You want to buy a new motorcycle. The payments will be $244.22 per month for 8 years. The interest rate is 7.35%. What is the asking price of the motorcycle?

OpenStudy (anonymous):

$244.22*=1953.76, 1953.76*.735=1436.0136, so 1953.76+1436.0136=$3389.7736, or $3389.77

OpenStudy (anonymous):

wouldnt it be 244.22*12=2930.64*8=23,445.12 times .735=17232.1632= 40,677

OpenStudy (anonymous):

yeah I totally left stuff out im sorry

OpenStudy (anonymous):

but yes that is the right way to do it I left out multiplying it by the 12 months sorry

OpenStudy (anonymous):

okay cool thx

OpenStudy (anonymous):

np